# The general position number under vertex and edge removal

### v1

@@ Abstract
Let ${\rm gp}(G)$ be the general position number of a graph $G$. It is proved that ${\rm gp}(G-x)\leq 2{\rm gp}(G)$ holds for any vertex $x$ of a connected graph $G$ and that if $x$ lies in some ${\rm gp}$-set of $G$, then ${\rm gp}(G)-1\leq{\rm gp}(G-x)$. Constructions are given which show that ${\rm gp}(G-x)$ can be much larger than ${\rm gp}(G)$ also when $G-x$ is connected. For diameter $2$ graphs it is proved that ${\rm gp}(G-x)\leq{\rm gp}(G)$, and that ${\rm gp}(G-x)\geq{\rm gp}(G)-1$ when the diameter of $G-x$ remains $2$. It is demonstrated that ${\rm gp}(G)/2\leq{\rm gp}(G-e)\leq 2{\rm gp}(G)$ holds for any edge $e$ of a graph $G$. For diameter $2$ graphs these results sharpens to ${\rm gp}(G)-1\leq{\rm gp}(G-e)\leq\ {\rm gp}(G)+1$. All these bounds are proved to be sharp.
@@ 1 Introduction
The general position problem was introduced to graph theory in [5, 17] and proved to be NP-hard in [17]. In [3], the structure of general position sets was clarified. Afterwards, the problem of determining the general position number of a graph received a wide attention, cf. [6, 23, 25, 28, 31]. This is in particular the case for graph products. The general position number of the Cartesian product of paths [13], of paths and cycles [16], and of two trees [30] were determined, while in [15] the general position number of strong product graphs was investigated. The concept has been modified and/or generalized into several directions. Let us point here to $d$-position sets [12], Steiner general position problem [11], edge general position problem [14, 19, 27], monophonic general position problem [26], and general position polynomials [8].
In this paper, we focus on how much the general position number of a graph can be affected by removing a vertex or by removing an edge. In the next section we give definitions and recall known results needed. In Section 3, we prove that ${\rm gp}(G-x)\leq 2{\rm gp}(G)$ holds for any vertex $x$ of a graph $G$. Then we demonstrate that ${\rm gp}(G-x)$ cannot be bounded from below by a function of ${\rm gp}(G)$. On the other hand, if $x$ lies in some ${\rm gp}$-set of $G$, then we prove that ${\rm gp}(G)-1\leq{\rm gp}(G-x)$. In Section 4, we give two constructions which show that ${\rm gp}(G-x)$ can be much larger than ${\rm gp}(G)$ also when $G-x$ is connected. In Section 5, we focus on the vertex removing operation in diameter $2$ graphs. We show that if ${\rm diam}(G)=2$, then ${\rm gp}(G-x)\leq{\rm gp}(G)$ and prove that ${\rm gp}(G)-1\leq{\rm gp}(G-x)\leq{\rm gp}(G)$ when the diameter of $G-x$ remains $2$. In Section 6, we prove that ${\rm gp}(G)/2\leq{\rm gp}(G-e)\leq 2{\rm gp}(G)$ holds for any edge $e$ of a graph $G$. For diameter $2$ graphs $G$ we sharpen the bound by proving that ${\rm gp}(G)-1\leq{\rm gp}(G-e)\leq\ {\rm gp}(G)+1$. All the above bounds are along the way shown to be sharp.
@@ 2 Preliminaries
The distance $d_{G}(u,v)$ between vertices $u$ and $v$ of $G$ is the number of edges on a shortest $u,v$-path. A shortest path of $G$ also is called a geodesic of $G$. The diameter of $G$ is the maximum distance between pairs of vertices of $G$ and is denoted by ${\rm diam}(G)$. A subgraph $H$ of $G$ is isometric if for each pair of vertices $u,v\in V(H)$ we have $d_{H}(u,v)=d_{G}(u,v)$. The interval between vertices $u$ and $v$ is
@@
A set $X\subseteq V(G)$ is a general position set of $G$ if for each pair $u,v\in X$ and any shortest $u,v$-path $P$ we have $V(P)\cap X=\{u,v\}$. The cardinality of a largest general position set of $G$ is the general position number of $G$ denoted by ${\rm gp}(G)$ and refereed to as the gp-number of $G$. A general position set $X$ of cardinality ${\rm gp}(G)$ is refereed to as a gp-set of $G$. For a vertex $u\in V(G)$, a set $X\subseteq V(G)$ is $u$-colinear if $X$ is a general position set such that $u\notin X$ and $y\notin I_{G}[x,u]$ for any $x,y\in X$.
@@
[29, Corollary 2.9] If $n\geq 4$, then ${\rm gp}(F_{n})=\lceil\frac{2n}{3}\rceil$.
The final known result we recall describes general position sets in an arbitrary graph. To state it, some more definitions are required. If ${\cal P}=\{S_{1},\ldots,S_{t}\}$ a partition of $S\subseteq V(G)$, then ${\cal P}$ is distance-constant if for any $i,j\in[t]$, $i\neq j$, there exists a constant $p_{ij}$, such that $d_{G}(x,y)=p_{ij}$ for every $x\in S_{i}$, $y\in S_{j}$. If so, we set $d_{G}(S_{i},S_{j})=p_{ij}$. A distance-constant partition ${\cal P}$ is in-transitive if $p_{ik}\neq p_{ij}+p_{jk}$ holds for $i,j,k\in[t]$.
@@ 3 General bounds
Proof. Let $R$ be an arbitrary ${\rm gp}$-set of $G-x$. Then clearly $x\not\in R$. We partition $R$ into two sets $R_{1}$ and $R_{2}$ as follows. For every $u\in R$, we put each vertex from $(I_{G}[u,x]\setminus\{u\})\cap R$ into $R_{2}$. This defines $R_{2}$, and then $R_{1}=R\setminus R_{2}$. By the definition of the sets $R_{1}$ and $R_{2}$, the set $R_{1}$ is an $x$-colinear set of $G$. We next claim that $R_{2}$ is also an $x$-colinear set of $G$.
Consider two vertices $u,u^{\prime}\in R_{2}$ and suppose by way of contradiction that there exists a shortest $x,u$-path $P_{x,u}$ in $G$ that contains $u^{\prime}$. Then $d_{G}(u,x)=d_{G}(u,u^{\prime})+d_{G}(u^{\prime},x)$. Since $u\in R_{2}$, there exists a vertex $u^{\prime\prime}\in R_{1}$, and a shortest $x,u^{\prime\prime}$-path $P_{x,u^{\prime\prime}}$ of $G$ passing through $u$. We are going to show that $u^{\prime\prime}$, $u$, and $u^{\prime}$ lie on a common shortest path in $G-x$. If not, it follows that there exists a shortest $u^{\prime},u^{\prime\prime}$-path in $G-x$ such that $d_{G-x}(u^{\prime},u^{\prime\prime})<d_{G-x}(u^{\prime},u)+d_{G-x}(u,u^{\prime\prime})$. From our assumption, $x$ does not lie on any shortest $u^{\prime},u^{\prime\prime}$-path, any shortest $u^{\prime},u$-path, and any shortest $u,u^{\prime\prime}$-path in $G$. Therefore, $d_{G-x}(u^{\prime},u^{\prime\prime})=d_{G}(u^{\prime},u^{\prime\prime})$, $d_{G-x}(u,u^{\prime\prime})=d_{G}(u,u^{\prime\prime})$, and $d_{G-x}(u,u^{\prime})=d_{G}(u,u^{\prime})$. Then
$d_{G}(x,u^{\prime\prime})$ $=d_{G}(x,u)+d_{G}(u,u^{\prime\prime})$
$=d_{G}(x,u^{\prime})+d_{G}(u^{\prime},u)+d_{G}(u,u^{\prime\prime})$
$=d_{G}(x,u^{\prime})+d_{G-x}(u^{\prime},u)+d_{G-x}(u,u^{\prime\prime})$
$>d_{G}(x,u^{\prime})+d_{G-x}(u^{\prime},u^{\prime\prime})$
$=d_{G}(x,u^{\prime})+d_{G}(u^{\prime},u^{\prime\prime})$
$\geq d_{G}(x,u^{\prime\prime})\,,$
which is not possible. This contradiction implies that $d_{G-x}(u^{\prime},u^{\prime\prime})=d_{G-x}(u^{\prime},u)+d_{G-x}(u,u^{\prime\prime})$, hence $u$ lies on a shortest $u^{\prime},u^{\prime\prime}$-path of $G-x$. Since we have assumed that $\{u,u^{\prime},u^{\prime\prime}\}\subseteq R$ is a general position set of $G-x$, we have obtained a contradiction which implies that $R_{2}$ is an $x$-colinear set of $G$.
We now claim that $R_{1}$ and $R_{2}$ are general position sets of $G$. Suppose first that $R_{1}$ is not a general position sets of $G$. Then there exist vertices $u,v,w\in R_{1}$ and a shortest $u,w$-path $P$ in $G$ that passes through $v$. Since $R_{1}\subseteq R$ and $R$ is a general position set of $G-x$, the path $P$ must contain the vertex $x$. But then $v\notin R_{1}$, a contradiction. Suppose second that $R_{2}$ is not a general position sets of $G$. Then there exist vertices $u,v,w\in R_{2}$ and a shortest $u,w$-path $P$ in $G$ that passes through $v$. Again we see that the path $P$ must contain the vertex $x$. Assume without loss of generality that $x$ lies between $u$ and $v$ on $P$. Since $w\in R_{2}$, there exists a vertex $w^{\prime}\in R_{1}$, such that $w$ is on a shortest $w^{\prime},x$-path. But then $w^{\prime},w,v$ lie on a common geodesic, a contradiction with the assumption that $R$ is a general position set of $G-x$.
Since $R_{1}$ and $R_{2}$ are general position sets of $G$, we can conclude that
@@
There is no general lower bound on ${\rm gp}(G-x)$ in terms of ${\rm gp}(G)$. To demonstrate it, consider the fan graphs $F_{n}$, $n\geq 3$. By Proposition 2.4 we have ${\rm gp}(F_{n})=\lceil\frac{2n}{3}\rceil$. Since clearly ${\rm gp}(F_{n}-x)=2$, where $x$ is the vertex of $F_{n}$ of degree $n$, we see that ${\rm gp}(G-x)$ can indeed be arbitrary smaller than ${\rm gp}(G)$. The next result leads to many additional such examples.
@@ 4 Two constructions
In Theorem 3.1 we have proved that the bound ${\rm gp}(G-x)\leq 2{\rm gp}(G)$ is sharp. However, the sharpness examples were such that $G-x$ is not connected. In this section we give two constructions which show that ${\rm gp}(G-x)$ can be much larger than ${\rm gp}(G)$ also when $G-x$ is connected.
@@
Proof. Set $Y_{j}=\{y_{1},\ldots,y_{j}\}$ for $j\in\{n,n+1\}$. We claim that $Y_{n+1}\cup Z_{n}$ is a general position set of $H_{n}$. Consider any two vertices $u$ and $v$ from $Y_{n+1}\cup Z_{n}$. If $u,v\in Y_{n+1}$, then $d_{H_{n}}(u,v)=1$. If $u,v\in Z_{n}$, then since $H_{n}[\{x,x^{\prime}\}\cup Z_{n}]\cong K_{2,n}$, we have $d_{H_{n}}(u,v)=2$. Finally, if $u\in Y_{n+1}$ and $v\in Z_{n}$, then $d_{H_{n}}(u,v)=2$. Hence, having Theorem 2.5 in mind, $Y_{n+1}\cup Z_{n}$ is a general position set of $H_{n}$. It follows that ${\rm gp}(H_{n})\geq|Y_{n+1}\cup Z_{n}|=2n+1$.
@@
We next give another family of graphs in which the general position number increases arbitrary by removing a vertex. If $k\geq 2$, then let the graph $G_{k}$ be constructed as follows. Let $V(G_{k})=X_{k}\cup Y_{k}\cup Z_{k+1}\cup\{w\}$, where $X_{k}=\{x_{1},\ldots,x_{k}\}$, $Y_{k}=\{y_{1},\ldots,y_{k}\}$, and $Z_{k+1}=\{z_{1},\ldots,z_{k+1}\}$. The vertex $w$ is adjacent to every vertex of $X_{k}\cup Y_{k}$, the vertices of $Z_{k+1}$ induce a complete graph $K_{k+1}$, $z_{2}$ is adjacent to $y_{2},\ldots,y_{k}$, and $z_{1}$ is adjacent to $y_{1}$, see Fig. 2. Note that ${\rm diam}(G_{k})=4$, and that $X_{k}$ and $Y_{k}$ are independent sets of vertices.
@@
Consider $G_{k}-z_{2}$. It is straightforward to check that $S=V(G_{k}-z_{2})\setminus\{w,y_{1},z_{1}\}$ is a general position set of $G_{k}-z_{2}$. For instance, any shortest path between a vertex $x_{i}\in X_{k}$ and any other vertex from $S$ avoids other vertices of $S$ because $d_{G_{k}-z_{2}}(x_{i},x_{j})=2$, $d_{G_{k}-z_{2}}(x_{i},y_{j})=2$, and $d_{G_{k}-z_{2}}(x_{i},z_{j})=4$ for $j\geq 3$. In the latter case, a shortest $x_{i},z_{j}$-path, induced by the vertices $x_{i},w,y_{1},z_{1}$, and $z_{j}$, is unique. Thus ${\rm gp}(G_{k}-z_{2})\geq 3k-2$. On the other hand, suppose that $S$ is a general position set of $G_{k}-z_{2}$ of size $13k-$ 1. Because $n(G_{k}-z_{2})=3k+1$, we have $S\cap\{w,y_{1},z_{1}\}\neq\emptyset$. But then we find a vertex from $X_{k}$, a vertex from $\{w,y_{1},z_{1}\}$, and a vertex from $Z_{k+1}\setminus\{z_{1},z_{2}\}$, such that these three vertices from $S$ and lie on a common shortest path in $G_{k}-z_{2}$. As a consequence, we conclude that ${\rm gp}(G_{k}-z_{2})\leq 3k-2$ and thus we have ${\rm gp}(G_{k}-z_{2})=3k-2$. $\square$
@@ 5 Vertex removing in diameter $2$ graphs
Note that for the graphs $H_{n}$ from Proposition 4.1 we have ${\rm diam}(H_{n})=3$ as well as ${\rm diam}(H_{n}-x)=3$. Also, for the graphs $G_{k}$ from Proposition 4.2 we have ${\rm diam}(G_{k})=4$ as well as ${\rm diam}(G_{k}-z_{2})=4$. In both cases we have seen that removing a vertex increases the general position number arbitrarily. In this section we therefore focus on the vertex removing operation in diameter $2$ graphs. We first show that in this case Theorem 3.1 sharpens as ${\rm gp}(G-x)\leq{\rm gp}(G)$. Second, we prove that if the diameter of $G-x$ remains $2$, then ${\rm gp}(G)-1\leq{\rm gp}(G-x)\leq{\rm gp}(G)$. Before presenting these results, we consider some examples.
@@
Another family which demonstrates sharpness of the upper bound in Theorem 5.2 is the family $K_{n}\boxtimes C_{m}$, where $n\geq 2$ and $m\in\{4,5\}$. By [15, Proposition 4.3] we have ${\rm gp}(K_{n}\boxtimes C_{4})=2n$ and ${\rm gp}(K_{n}\boxtimes C_{5})=3n$. Moreover, it is straightforward to check that the general position number does not change after one vertex is removed from these graphs.

### v2

@@ Abstract
Let ${\rm gp}(G)$ be the general position number of a graph $G$. It is proved that ${\rm gp}(G-x)\leq 2{\rm gp}(G)$ holds for any vertex $x$ of a connected graph $G$ and that if $x$ lies in some ${\rm gp}$-set of $G$, then ${\rm gp}(G)-1\leq{\rm gp}(G-x)$. Constructions are given which show that ${\rm gp}(G-x)$ can be much larger than ${\rm gp}(G)$ also when $G-x$ is connected. For diameter $2$ graphs it is proved that ${\rm gp}(G-x)\leq{\rm gp}(G)$, and that ${\rm gp}(G-x)\geq{\rm gp}(G)-1$ when the diameter of $G-x$ remains $2$. It is demonstrated that ${\rm gp}(G)/2\leq{\rm gp}(G-e)\leq 2{\rm gp}(G)$ holds for any edge $e$ of a graph $G$. For diameter $2$ graphs these results can be improved to ${\rm gp}(G)-1\leq{\rm gp}(G-e)\leq\ {\rm gp}(G)+1$. All these bounds are proved to be sharp.
@@ 1 Introduction
The general position problem was introduced to graph theory in [5, 17] and proved to be NP-hard in [17]. In [3], the structure of general position sets was clarified. Afterwards, the problem of determining the general position number of a graph received wide attention, cf. [6, 23, 25, 26, 28, 30, 33]. This is particularly the case for graph products. The general position number of the Cartesian product of paths [13], of paths and cycles [16], and of two trees [32] were determined, while in [15] the general position number of strong product graphs was investigated. The concept has been modified and/or generalized into several directions. Let us point here to $d$-position sets [12], Steiner general position problem [11], edge general position problem [14, 19, 29], monophonic general position problem [27], and general position polynomials [8].
In this paper, we focus on how much the general position number of a graph can be affected by removing a vertex or by removing an edge. In the next section we give definitions and recall some known results that will be needed later on. In Section 3, we prove that ${\rm gp}(G-x)\leq 2{\rm gp}(G)$ holds for any vertex $x$ of a graph $G$. Then we demonstrate that ${\rm gp}(G-x)$ cannot be bounded from below by a function of ${\rm gp}(G)$. On the other hand, if $x$ lies in some ${\rm gp}$-set of $G$, then we prove that ${\rm gp}(G)-1\leq{\rm gp}(G-x)$. In Section 4, we give two constructions which show that ${\rm gp}(G-x)$ can be much larger than ${\rm gp}(G)$ also when $G-x$ is connected. In Section 5, we focus on the vertex removing operation in diameter $2$ graphs. We show that if ${\rm diam}(G)=2$, then ${\rm gp}(G-x)\leq{\rm gp}(G)$ and prove that ${\rm gp}(G)-1\leq{\rm gp}(G-x)\leq{\rm gp}(G)$ when the diameter of $G-x$ remains $2$. In Section 6, we prove that ${\rm gp}(G)/2\leq{\rm gp}(G-e)\leq 2{\rm gp}(G)$ holds for any edge $e$ of a graph $G$. For diameter $2$ graphs $G$ we sharpen the bound by proving that ${\rm gp}(G)-1\leq{\rm gp}(G-e)\leq\ {\rm gp}(G)+1$. All the above bounds are along the way shown to be sharp. We conclude the paper by listing several directions for future investigations.
@@ 2 Preliminaries
The distance $d_{G}(u,v)$ between vertices $u$ and $v$ of $G$ is the number of edges on a shortest $u,v$-path. A shortest path of $G$ is also called a geodesic of $G$. The diameter of $G$ is the maximum distance between pairs of vertices of $G$ and is denoted by ${\rm diam}(G)$. A subgraph $H$ of $G$ is isometric if for each pair of vertices $u,v\in V(H)$ we have $d_{H}(u,v)=d_{G}(u,v)$. The interval between vertices $u$ and $v$ is
@@
A set $X\subseteq V(G)$ is a general position set of $G$ if for each pair $u,v\in X$ and any shortest $u,v$-path $P$ we have $V(P)\cap X=\{u,v\}$. The cardinality of a largest general position set of $G$ is the general position number of $G$ denoted by ${\rm gp}(G)$ and referred to as the gp-number of $G$. A general position set $X$ of cardinality ${\rm gp}(G)$ is referred to as a gp-set of $G$.
@@
[31, Corollary 2.9] If $n\geq 4$, then ${\rm gp}(F_{n})=\lceil\frac{2(n+1)}{3}\rceil$.
The final known result we recall describes general position sets in an arbitrary graph. To state it, some more definitions are required. If ${\cal P}=\{S_{1},\ldots,S_{t}\}$ is a partition of $S\subseteq V(G)$, then ${\cal P}$ is distance-constant if for any $i,j\in[t]$, $i\neq j$, there exists a constant $p_{ij}$, such that $d_{G}(x,y)=p_{ij}$ for every $x\in S_{i}$, $y\in S_{j}$. If so, we set $d_{G}(S_{i},S_{j})=p_{ij}$. A distance-constant partition ${\cal P}$ is in-transitive if $p_{ik}\neq p_{ij}+p_{jk}$ holds for $i,j,k\in[t]$.
@@ 3 General bounds
Proof. Let $R$ be an arbitrary ${\rm gp}$-set of $G-x$. Then clearly $x\not\in R$. We partition $R$ into two sets $R_{1}$ and $R_{2}$ as follows. For every $u\in R$, we put each vertex from $(I_{G}[u,x]\setminus\{u\})\cap R$ into $R_{2}$. This defines $R_{2}$, and then $R_{1}=R\setminus R_{2}$. We claim that $R_{1}$ and $R_{2}$ are general position sets of $G$.
Suppose first that $R_{1}$ is not a general position set of $G$. Then there exist vertices $u,v,w\in R_{1}$ and a shortest $u,w$-path $P$ in $G$ that passes through $v$. Since $R_{1}\subseteq R$ and $R$ is a general position set of $G-x$, the path $P$ must contain the vertex $x$. We may without loss of generality assume that $x$ lies in the $u,v$-subpath of $P$. Then the $w,x$-subpath of $P$ is a shortest $w,x$-path in $G$. By definition of $R_{2}$ we get that $v\in R_{2}$, a contradiction.
Suppose second that $R_{2}$ is not a general position set of $G$. Hence there exist vertices $u,v,w\in R_{2}$ and a shortest $u,w$-path $P$ in $G$ that passes through $v$. Just as in the above paragraph, the path $P$ must contain the vertex $x$ and we may assume that $x$ lies in the $u,v$-subpath of $P$. Since $w\in R_{2}$, there exists a vertex $w^{\prime}\in R$, such that $w$ lies on a shortest $w^{\prime},x$-path $Q$ in $G$. Since $P$ and $Q$ are shortest paths, we infer that the $w,x$-subpath of $Q$ and the $w,x$-subpath of $P$ are of the same length. But this in turn implies that the vertices $w^{\prime},w,v$ from $R$ lie on a common shortest path in $G-x$, a contradiction.
We have thus proved that $R_{1}$ and $R_{2}$ are general position sets of $G$. Therefore,
@@
There is no general lower bound on ${\rm gp}(G-x)$ in terms of ${\rm gp}(G)$. To demonstrate it, consider the fan graphs $F_{n}$, $n\geq 3$. By Proposition 2.4 we have ${\rm gp}(F_{n})=\lceil\frac{2(n+1)}{3}\rceil$. Since clearly ${\rm gp}(F_{n}-x)=2$, where $x$ is the vertex of $F_{n}$ of degree $n$, we see that ${\rm gp}(G-x)$ can indeed be arbitrarily smaller than ${\rm gp}(G)$. The next result leads to many additional such examples.
@@ 4 Two constructions
In Theorem 3.1 we have proved that the bound ${\rm gp}(G-x)\leq 2{\rm gp}(G)$ is sharp. However, sharpness examples were such that $G-x$ is not connected. In this section we give two constructions which show that ${\rm gp}(G-x)$ can be much larger than ${\rm gp}(G)$ also when $G-x$ is connected.
@@
Proof. Set $Y_{j}=\{y_{1},\ldots,y_{j}\}$ for $j\in\{n,n+1,2n\}$. Using Theorem 2.5, we see that $Y_{n+1}\cup Z_{n}$ is a general position set of $H_{n}$. It follows that ${\rm gp}(H_{n})\geq|Y_{n+1}\cup Z_{n}|=2n+1$.
@@
We next give another family of graphs in which the general position number increases arbitrarily by removing a vertex. If $k\geq 2$, then let the graph $G_{k}$ be constructed as follows. Let $V(G_{k})=X_{k}\cup Y_{k}\cup Z_{k+1}\cup\{w\}$, where $X_{k}=\{x_{1},\ldots,x_{k}\}$, $Y_{k}=\{y_{1},\ldots,y_{k}\}$, and $Z_{k+1}=\{z_{1},\ldots,z_{k+1}\}$. The vertex $w$ is adjacent to every vertex of $X_{k}\cup Y_{k}$, the vertices of $Z_{k+1}$ induce a complete graph $K_{k+1}$, $z_{2}$ is adjacent to $y_{2},\ldots,y_{k}$, and $z_{1}$ is adjacent to $y_{1}$, see Fig. 2. Note that ${\rm diam}(G_{k})=4$, and that $X_{k}$ and $Y_{k}$ are independent sets of vertices.
@@
Consider $G_{k}-z_{2}$. It is straightforward to check that $S=V(G_{k}-z_{2})\setminus\{w,y_{1},z_{1}\}$ is a general position set of $G_{k}-z_{2}$. For instance, any shortest path between a vertex $x_{i}\in X_{k}$ and any other vertex from $S$ avoids other vertices of $S$ because $d_{G_{k}-z_{2}}(x_{i},x_{j})=2$, $d_{G_{k}-z_{2}}(x_{i},y_{j})=2$, and $d_{G_{k}-z_{2}}(x_{i},z_{j})=4$ for $j\geq 3$. In the latter case, a shortest $x_{i},z_{j}$-path, induced by the vertices $x_{i},w,y_{1},z_{1}$, and $z_{j}$, is unique. Thus ${\rm gp}(G_{k}-z_{2})\geq 3k-2$. On the other hand, suppose that $S$ is a general position set of $G_{k}-z_{2}$ of size $13k-$ 1. Because $n(G_{k}-z_{2})=3k+1$, we have $S\cap\{w,y_{1},z_{1}\}\neq\emptyset$. But then we find a vertex from $X_{k}$, a vertex from $\{w,y_{1},z_{1}\}$, and a vertex from $Z_{k+1}\setminus\{z_{1},z_{2}\}$, such that these three vertices from $S$ lie on a common shortest path in $G_{k}-z_{2}$. As a consequence, we conclude that ${\rm gp}(G_{k}-z_{2})\leq 3k-2$ and thus we have ${\rm gp}(G_{k}-z_{2})=3k-2$. $\square$
@@ 5 Vertex removing in diameter $2$ graphs
Note that for the graphs $H_{n}$ from Proposition 4.1 we have ${\rm diam}(H_{n})=3$ as well as ${\rm diam}(H_{n}-x)=3$. Also, for the graphs $G_{k}$ from Proposition 4.2 we have ${\rm diam}(G_{k})=4$ as well as ${\rm diam}(G_{k}-z_{2})=4$. In both cases we have seen that removing a vertex increases the general position number arbitrarily. In this section we therefore focus on the vertex removing operation in diameter $2$ graphs. We first show that in this case Theorem 3.1 can be sharpened to ${\rm gp}(G-x)\leq{\rm gp}(G)$. Second, we prove that if the diameter of $G-x$ remains $2$, then ${\rm gp}(G)-1\leq{\rm gp}(G-x)\leq{\rm gp}(G)$. Before presenting these results, we consider some examples.
@@
We note that the bounds from Proposition 5.1 and Theorem 5.2 can also be deduced from the fact that a subset of a diameter two graph is in general position if and only if it is an independent union of cliques.
Another family which demonstrates sharpness of the upper bound in Theorem 5.2 is the family of strong products $K_{n}\boxtimes C_{m}$, where $n\geq 2$ and $m\in\{4,5\}$. By [15, Proposition 4.3] we have ${\rm gp}(K_{n}\boxtimes C_{4})=2n$ and ${\rm gp}(K_{n}\boxtimes C_{5})=3n$. Moreover, it is straightforward to check that the general position number does not change after one vertex is removed from these graphs.